carmelle:

Factor x^2 − 2x + 3. (x − 3)(x − 1) (x + 3)(x + 1) (x − 3)(x + 1) Prime

AZ:

what two numbers multiply to give you 3 and add up to -2

AZ:

since 3 is a positive number and it has to add up to a negative number that means both the numbers have to be negative

carmelle:

uh, idk

carmelle:

wait.. -3 x -1

jhonyy9:

x^2 -2x +3 = 0 this quadratic has the discriminant smaller then zero so has non real roots

carmelle:

im confused ._.

carmelle:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 x^2 -2x +3 = 0 this quadratic has the discriminant smaller then zero so has non real roots \(\color{#0cbb34}{\text{End of Quote}}\) so its prime?

carmelle:

what's wrong @snowflake0531 ?

jhonyy9:

@AZ do you see it now ?

snowflake0531:

Like, you can't factor it out the regular way

snowflake0531:

I just saw it too lol.

carmelle:

oh :/

carmelle:

ok, so its prime... thanks for the help :>

carmelle:

ok

AZ:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 x^2 -2x +3 = 0 this quadratic has the discriminant smaller then zero so has non real roots \(\color{#0cbb34}{\text{End of Quote}}\) ahh, you are correct

AZ:

so what Jhonyy is saying- you can use the discriminant to tell you about the roots the discriminant is \( b^2 - 4ac\) when the equation is written as \( ax^2 + bx + c = 0\)

1 attachment
carmelle:

ah, ok I get it now

AZ:

so (-2)^2 - 4(1)(3) = 4 - 12 = -8 since it's negative, we would have imaginary roots

carmelle:

okay I get it

AZ:

Good :) Thank you @jhonyy9 for pointing that out, I completely missed it :))

AZ:

you would have also realized, there are no 2 numbers that can multiply to give you 3 and add up to -2 -3 and -1 would add up to -4 not -2

carmelle:

oh yeah

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @AZ Good :) Thank you @jhonyy9 for pointing that out, I completely missed it :)) \(\color{#0cbb34}{\text{End of Quote}}\) np